OpenStudy (anonymous):

Angle 1 and Angle 2 are both supplementary angles. m angle 1 is 4y + 7, and m angle 2 is 9y + . Find m angle 2. (2 points) please show step-by-step

OpenStudy (anonymous):

Step #1: Supplementary angles add up to be 180 degrees. Step #2: Let's look at the angles: Angle 1: 4y + 7 Angle 2: 9y (4y + 7) + 9y = 180 13y + 7 = 180 13y = 173 y = 13.3 (Approximation) Step #3: Plug back into equation #1 and equation #2 Angle 1: 4y + 7 = 4*(13.3) + 7 = 60.2 degrees Angle 2: 9*y = 119.7 degrees I hope this helps =D I have a hunch that there is something wrong with how the problem is worded when it is typed "Angle 2 is equal to 9y + ". Could there be some holes in how the problem was worded?

OpenStudy (anonymous):

9y + 4!! sorry haha

OpenStudy (anonymous):

Could you possibly redo angle 2? Thank you for the step-by-step!!

OpenStudy (anonymous):

I will also give you a a medal for helping me of course (:

OpenStudy (anonymous):

Here is what I have below. (4y + 7) + (9y + 4) = 180 13y + 11 = 180 13y = 169 y = 13. Angle 1: (4*13)+7 = 59 degrees Angle 2: (9*(13)) + 4 = 121 degrees I hope this helps =D. If 121 degrees is added to 59 degrees, the answer is 180 degrees.

OpenStudy (anonymous):

Thank you so much, it wouldn't be to much to ask for some more help? you are really great!

OpenStudy (anonymous):

And if you had to explain each step, how would you?